Jimbo will receive $X every 8 months, starting 10 months from today. He will receive these payments forever. His stated annual interest is 6%, compounded every two months. Find X if the present value of all of the payments is $75,000.

I know the answers is c=$3,075.75 but I don't know how to solve it

Can someone help me solve this without using excel?
Answer #1

PV = $75000 , ANNUAL INTEREST = 6% , receive $X every 8 month .

8 - month discount rate = 1.014 - 1 = 0.04060401

$75000 = ( x / 0.04060401 ) ( 1 / 1.01 )

$75000 = 24.3843 c

c = 75000 / 24.3843

therefor , c = $3075.75